# Scheduling problem of aircrafts on a same runway and dual runways

### v5

@@ VI Scheduling problem of landing and takeoff aircraft on dual runways
(1) Suppose that $Tcf_{i_{1}-1}\in\phi_{0}$, $Tcf_{i_{1}}\in\phi_{1}$, for the sequence $\phi_{0}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{1}$, there is a path from its second last aircraft to its first aircraft. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a T-block subsequence of $\Phi$, and $(|\phi_{1}|-1)/(|\phi_{0}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the takeoff time increment of T-block subsequence, where $Tcf_{i_{2}}$ is the second last takeoff aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the takeoff sequence in the T-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the landing times of the landing aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and the takeoff times of the takeoff aircraft satisfy that $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, no conflicts of separation times would occur in $\bar{\phi}$ and if $\mu>\mu_{0}$, conflicts of separation times would occur in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the T-block subsequence $\phi$.
(2) Suppose that $Tcf_{i_{1}-1}\in\phi_{1}$, $Tcf_{i_{1}}\in\phi_{0}$, for the sequence $\phi_{1}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{0}$, there is a path from the second last aircraft to the first aircraft in $\phi_{0}$. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a D-block subsequence of $\Phi$, and $(|\phi_{0}|-1)/(|\phi_{1}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the landing time increment of D-block subsequence, where $Tcf_{i_{2}}$ is the second last landing aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the landing sequence in the D-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the takeoff times of the takeoff aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, no conflicts of separation times would occur in $\bar{\phi}$ and if $\mu>\mu_{0}$, conflicts of separation times would occur in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the D-block subsequence $\phi$.
@@
In Tables III and IV, the separation times between aircraft in the landing subsequence $\phi_{0}=\langle Tcf_{1}^{0},Tcf_{2}^{0}\rangle$ and the landing time increments of D-block subsequences are shown, where the rows represent the takeoff separation time vector (TSTV in short) between aircraft in the takeoff subsequence $\phi_{1}=\langle Tcf^{1}_{1},Tcf^{1}_{2},Tcf^{1}_{3},Tcf^{1}_{4}\rangle$, and the columns represent the minimum landing separation time (MLST) between aircraft in $\phi_{0}$.
@@
From Tables I-IV, it can be seen that once the separation times between the landing/takeoff aircraft are fixed, in most of the scenarios, each of the takeoff/landing aircraft is relevant to an aircraft with different operation task, which form block subsequences of length $1/(*)$. There are only a very small number of scenarios marked in red such that each of the takeoff/landing aircraft is only relevant to an aircraft with the same operation task and block subsequences are not formed. For example, consider an optimal matching problem between two sequences, $\phi_{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3}\rangle$ and $\phi_{1}=\langle Tcf_{4},Tcf_{5},Tcf_{6}\rangle$, where $S_{1}(\phi_{0})=S_{4}(\phi_{1})$, $Y_{12}=Y_{23}=90$, $Y_{45}=60$, and $Y_{56}=80$. If we let $S_{12}(\phi_{0})=S_{23}(\phi_{0})=90$, $S_{45}(\phi_{1})$ and $S_{56}(\phi_{1})$ should be taken as $S_{45}(\phi_{1})=S_{56}(\phi_{1})=90$, whereas if we let $S_{45}(\phi_{1})=60$ and $S_{56}(\phi_{1})=80$, $S_{12}(\phi_{0})$ should be taken as $S_{12}(\phi_{0})=140$, and $S_{23}(\phi_{0})$ can be taken as $S_{23}(\phi_{0})=90$ if there are no other aircraft. In this example, it can be seen that the matching between a landing sequence and a takeoff sequence is not very complicated and the number of all possible block subsequences is $2$.
@@
From Tables V-VIII, it can be seen that the range of the time increments of block subsequences is not large as well, in particular, when the takeoff time increments are considered. As a matter of fact, based on the calculations for more general scenarios including those listed in Tables I-VIII, it is found that if the number of breakpoint aircraft in a block subsequence is fixed, the number of all possible block subsequences of different class combinations and the range of the takeoff/landing time increments of block subsequences are both not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system in the absence of resident-point aircraft. This property can be fully explored to find the optimal sequence when a landing sequence and a takeoff sequence are matched.
@@ VII-A Some necessary results
It can be easily seen that the minimum insertion/extraction time increment is only related to the aircraft classes. Since the number of the aircraft classes is $\eta$, all possible minimum insertion/extraction time increments can easily be calculated out.
@@
For the optimization problem (6), the set of the optimal sequences might be composed of the breakpoint-drift equivalent transformation sets spanned by multiple optimal sequence basis. For example, consider a sequence $\phi_{a}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ and $\phi_{b}=\langle Tcf_{1},Tcf_{3},Tcf_{4},Tcf_{2},Tcf_{5}\rangle$, where $Y_{12}=180,Y_{13}=160,Y_{23}=Y_{25}=Y_{34}=60$, $Y_{45}=120$, and $Y_{42}=140$. It is clear that $F(\phi_{a})=F(\phi_{b})$ but the separation times between the breakpoint aircraft and their trailing aircraft in $\phi_{a}$ and $\phi_{b}$ are different.
@@
It should be noted that this lemma can be used in the proposed algorithms to easily obtain the minimum and maximum adjustment ranges of all aircraft, which can be applied to find all breakpoint-drift equivalent transformations of an optimal sequence. For example, based on $\Phi_{a}$ in Lemma 17, generate a sequence $\Phi_{b1}$ containing $\langle Tcf_{i},Tcf_{i+3}\rangle$ as its subsequence by moving the aircraft $Tcf_{i+1},Tcf_{i+2}$ to be between $Tcf_{k_{0}}$ and $Tcf_{k_{0}+1}$ and between $Tcf_{k_{1}}$ and $Tcf_{k_{1}+1}$ for $k_{0}<k_{1}<i$. From Lemma 17, we can first perform the movement of the aircraft $Tcf_{i+1}$ and then the movement the aircraft $Tcf_{i+2}$. If we only perform the movement of the aircraft $Tcf_{i+2}$, the generated sequence $\Phi_{c}$ might not exist because a possible case satisfying that $S_{i+2}(\Phi_{c})\notin[f^{\min}_{i+2},f^{\max}_{i+2}]$ and $S_{i+2}(\Phi_{b1})\in[f^{\min}_{i+2},f^{\max}_{i+2}]$ might occur.
In the following, we give an example to show how to obtain the maximum number of aircraft that can be moved to be before one aircraft based on Lemma 17 and the analysis of the above example. Consider a sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{s_{0}},Tcf_{s_{0}+1},\cdots,Tcf_{n}\rangle$, where $\phi_{1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{s_{0}}\rangle$ is a class-monotonically-decreasing sequence. The objective is to move the aircraft $Tcf_{s_{0}+1},\cdots,Tcf_{n}$ as many as possible to be before $Tcf_{s_{0}}$ without generating new breakpoint aircraft in the generated subsequence $\langle Tcf_{a},\cdots,Tcf_{s_{0}}\rangle$. To this end, we first rearrange the aircraft $Tcf_{s_{0}+1},\cdots,Tcf_{n}$ in descending order of their aircraft classes, where the aircraft of the same class are rearranged in ascending order of their earliest landing/takeoff times. Let $\phi_{2}=\langle Tcf^{a}_{s_{0}+1},\cdots,Tcf^{a}_{n}\rangle$ denote the generated sequence of the aircraft $Tcf_{s_{0}+1},\cdots,Tcf_{n}$. Attempt to insert the aircraft in the order of $\phi_{2}$ into the subsequence $\phi_{1}$ and we can find the maximum number of aircraft that can be moved to be before aircraft $Tcf_{s_{0}}$.
@@
In the following theorem, we make an estimation about the number of breakpoint aircraft for landing/takeoff sequence.
Theorem 18
Consider a landing/takeoff sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Suppose that $[f^{0},f^{1}]\subset\bigcap_{i=1}^{n}[f_{i}^{\min},f_{i}^{\max}]$. In the time interval $[f^{0},f^{1}]$, the minimum number of breakpoint aircraft lie in $[0,2]$.
Proof: Since $[f^{0},f^{1}]\subset\bigcap_{i=1}^{n}[f_{i}^{\min},f_{i}^{\max}]$, the orders of aircraft in the time interval $[f^{0},f^{1}]$ can be adjusted arbitrarily. Thus, there might be 00 breakpoint aircraft. Note that the aircraft in the time interval $[t_{0},f^{0})$ and the aircraft in the time interval $[f^{1},+\infty)$ might form class-monotonically-decreasing sequences that finally generate $1$ or $2$ breakpoint aircraft in the time interval $[f^{0},f^{1}]$. Summarizing all the analysis, in the time interval $[f^{0},f^{1}]$, the minimum number of breakpoint aircraft lie in $[0,2]$.

### v7

@@ VI Scheduling problem of landing and takeoff aircraft on dual runways
(1) Suppose that $Tcf_{i_{1}-1}\in\phi_{0}$, $Tcf_{i_{1}}\in\phi_{1}$, for the sequence $\phi_{0}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{1}$, there is a path from its second last aircraft to its first aircraft. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a T-block subsequence of $\Phi$, and $(|\phi_{1}|-1)/(|\phi_{0}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the takeoff time increment of T-block subsequence, where $Tcf_{i_{2}}$ is the second last takeoff aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the takeoff sequence in the T-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the landing times of the landing aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and the takeoff times of the takeoff aircraft satisfy that $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, the minimum separation requirements can be satisfied in $\bar{\phi}$ and if $\mu>\mu_{0}$, the minimum separation requirements cannot be satisfied in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the T-block subsequence $\phi$.
(2) Suppose that $Tcf_{i_{1}-1}\in\phi_{1}$, $Tcf_{i_{1}}\in\phi_{0}$, for the sequence $\phi_{1}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{0}$, there is a path from the second last aircraft to the first aircraft in $\phi_{0}$. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a D-block subsequence of $\Phi$, and $(|\phi_{0}|-1)/(|\phi_{1}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the landing time increment of D-block subsequence, where $Tcf_{i_{2}}$ is the second last landing aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the landing sequence in the D-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the takeoff times of the takeoff aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, the minimum separation requirements can be satisfied in $\bar{\phi}$ and if $\mu>\mu_{0}$, the minimum separation requirements cannot be satisfied in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the D-block subsequence $\phi$.
@@
Note that in the definitions of T-block subsequences and D-subsequences, the last aircraft $Tcf_{i_{1}}$ is a semi-resident-point aircraft.
@@
In Tables III, IV, VII and VIII, the separation times between aircraft in the landing subsequence $\phi_{0}=\langle Tcf_{1}^{0},Tcf_{2}^{0}\rangle$ and the landing time increments of D-block subsequences are shown, where the rows represent the takeoff separation time vector (TSTV in short) between aircraft in the takeoff subsequence $\phi_{1}=\langle Tcf^{1}_{1},Tcf^{1}_{2},Tcf^{1}_{3},Tcf^{1}_{4}\rangle$, and the columns represent the minimum landing separation time (MLST) between aircraft in $\phi_{0}$.
@@
From Tables I-IV, it can be seen that once the separation times between the landing/takeoff aircraft are fixed, in most of the scenarios, each of the takeoff/landing aircraft is relevant to an aircraft with different operation task, which form block subsequences of length $1/(*)$. There are only a very small number of scenarios marked in red such that each of the takeoff/landing aircraft is relevant to only an aircraft with the same operation task and block subsequences are not formed. For example, consider an optimal matching problem between two sequences, $\phi_{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3}\rangle$ and $\phi_{1}=\langle Tcf_{4},Tcf_{5},Tcf_{6}\rangle$, where $S_{1}(\phi_{0})=S_{4}(\phi_{1})$, $Y_{12}=Y_{23}=90$, $Y_{45}=60$, and $Y_{56}=80$. If we let $S_{12}(\phi_{0})=S_{23}(\phi_{0})=90$, $S_{45}(\phi_{1})$ and $S_{56}(\phi_{1})$ should be taken as $S_{45}(\phi_{1})=S_{56}(\phi_{1})=90$, whereas if we let $S_{45}(\phi_{1})=60$ and $S_{56}(\phi_{1})=80$, $S_{12}(\phi_{0})$ should be taken as $S_{12}(\phi_{0})=140$, and $S_{23}(\phi_{0})$ can be taken as $S_{23}(\phi_{0})=90$ if there are no other aircraft. In this example, it can be seen that the matching between a landing sequence and a takeoff sequence is not very complicated and the number of all possible block subsequences is $2$.
From Tables V-VIII, it can be seen that the range of the time increments of block subsequences is not large as well, in particular when the takeoff time increments are considered. As a matter of fact, based on the calculations for more general scenarios including those listed in Tables I-VIII, it is found that if the number of breakpoint aircraft in a block subsequence is fixed, the number of all possible block subsequences of different class combinations and the range of the takeoff/landing time increments of block subsequences are both not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system in the absence of resident-point aircraft. This property can be fully explored to find the optimal sequence when a landing sequence and a takeoff sequence are matched.
@@ VII-A Some necessary results
It can be easily seen that the minimum insertion/extraction time increment is related to only the aircraft classes. Since the number of the aircraft classes is $\eta$, all possible minimum insertion/extraction time increments can easily be calculated out.
@@
For the optimization problem (6), the set of the optimal sequences might be composed of the breakpoint-drift equivalent transformation sets spanned by multiple optimal sequence basis. For example, consider a landing sequence $\phi_{a}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ and $\phi_{b}=\langle Tcf_{1},Tcf_{3},Tcf_{4},Tcf_{2},Tcf_{5}\rangle$ under Assumptions 1-5, where $Y_{12}=180,Y_{13}=160,Y_{23}=Y_{25}=Y_{34}=60$, $Y_{45}=120$, and $Y_{42}=140$. It is clear that $F(\phi_{a})=F(\phi_{b})$ but the separation times between the breakpoint aircraft and their trailing aircraft in $\phi_{a}$ and $\phi_{b}$ are different.
@@
Lemma 17 shows the impact of aircraft order adjustments on the other aircraft. This lemma can be used to easily obtain the minimum and maximum adjustment ranges of aircraft by first making adjustments that have little impact on other aircrafts, which can be applied to find all breakpoint-drift equivalent transformations of an optimal sequence. For example, based on $\Phi_{a}$ in Lemma 17, suppose that a sequence $\Phi_{b1}$ containing a subsequence $\langle Tcf_{i},Tcf_{i+3}\rangle$ can be generated by moving the aircraft $Tcf_{i+1},Tcf_{i+2}$ to be between $Tcf_{k_{0}}$ and $Tcf_{k_{0}+1}$ and between $Tcf_{k_{1}}$ and $Tcf_{k_{1}+1}$ for $k_{0}<k_{1}<i$. From Lemma 17, we can first move $Tcf_{i+1}$ and then move $Tcf_{i+2}$ since the movement of $Tcf_{i+1}$ has less effects on $Tcf_{i+2}$. If we first perform the movement of the aircraft $Tcf_{i+2}$ and keep the order of $Tcf_{i+1}$ unchanged, the generated sequence $\Phi_{c}$ might not exist because a possible case satisfying that $S_{i+2}(\Phi_{c})\notin[f^{\min}_{i+2},f^{\max}_{i+2}]$ and $S_{i+2}(\Phi_{b1})\in[f^{\min}_{i+2},f^{\max}_{i+2}]$ might occur.
In the following, we give an example to show how to obtain the maximum number of aircraft that can be moved to be before one aircraft based on Lemma 17 and the analysis of the above example. Consider a sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{s_{0}},Tcf_{s_{0}+1},\cdots,Tcf_{n}\rangle$, where $\phi_{1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{s_{0}}\rangle$ is a class-monotonically-decreasing sequence. The objective is to move the aircraft $Tcf_{s_{0}+1},\cdots,Tcf_{n}$ as many as possible to be before $Tcf_{s_{0}}$ without generating new breakpoint aircraft. To this end, we first rearrange the aircraft $Tcf_{s_{0}+1},\cdots,Tcf_{n}$ in descending order of their aircraft classes, where the aircraft of the same class are rearranged in ascending order of their earliest landing/takeoff times. Let $\phi_{2}=\langle Tcf^{a}_{s_{0}+1},\cdots,Tcf^{a}_{n}\rangle$ denote the generated sequence of the aircraft $Tcf_{s_{0}+1},\cdots,Tcf_{n}$. Attempt to insert the aircraft in the order of $\phi_{2}$ into the subsequence $\phi_{1}$ and we can roughly find the maximum number of aircraft that can be moved to be before aircraft $Tcf_{s_{0}}$.
@@
In the following proposition, we make an estimation about the number of breakpoint aircraft for a landing/takeoff sequence.
Proposition 5
Consider a landing/takeoff sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Suppose that $(\phi,Sr(\phi))$ is an optimal solution of the optimization problem (6) for aircraft $Tcf_{1},Tcf_{2},\cdots,Tcf_{n}$, and $[f^{0},f^{1}]\subset\bigcap_{i=1}^{n}[f_{i}^{\min},$ $f_{i}^{\max}]$. In the time interval $[f^{0},f^{1}]$, the number of breakpoint aircraft lies in $[0,2]$.
Proof: Since $[f^{0},f^{1}]\subset\bigcap_{i=1}^{n}[f_{i}^{\min},f_{i}^{\max}]$, the orders of aircraft in the time interval $[f^{0},f^{1}]$ can be adjusted arbitrarily. Thus, there might be 00 breakpoint aircraft. Note that the aircraft in the time interval $[t_{0},f^{0})$ and the aircraft in the time interval $[f^{1},+\infty)$ might form class-monotonically-decreasing sequences that finally generate $1$ or $2$ breakpoint aircraft in the time interval $[f^{0},f^{1}]$. Summarizing all the analysis, in the time interval $[f^{0},f^{1}]$, the number of breakpoint aircraft lies in $[0,2]$.
@@
Theorem 19
Consider a mixed landing and takeoff sequence $\Phi_{a}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4}\rangle$ on a same runway, where $S_{12}(\Phi_{a})\geq Y_{12}$, aircraft $Tcf_{3}$ is relevant to aircraft $Tcf_{1}$ or $Tcf_{2}$ and aircraft $Tcf_{4}$ is relevant to aircraft $Tcf_{2}$ or $Tcf_{3}$. Generate a new sequence $\Phi_{b}$ by inserting an aircraft $Tcf_{5}$ to be between aircraft $Tcf_{2}$ and $Tcf_{3}$, where $S_{1}(\Phi_{a})=S_{1}(\Phi_{b})$, $S_{2}(\Phi_{a})=S_{2}(\Phi_{b})$, aircraft $Tcf_{5}$ is relevant to aircraft $Tcf_{1}$ or $Tcf_{2}$, aircraft $Tcf_{3}$ is relevant to aircraft $Tcf_{2}$ or $Tcf_{5}$ and aircraft $Tcf_{4}$ is relevant to aircraft $Tcf_{5}$ or $Tcf_{3}$. Let $\omega_{min}(\Phi_{a},Tcf_{5})=\min\{S_{3}(\Phi_{b})-S_{3}(\Phi_{a}),S_{4}(\Phi_{b})-S_{4}(\Phi_{a})\}$. The following statements hold.
(1) Suppose that aircraft $Tcf_{2}$, $Tcf_{3}$ and $Tcf_{5}$ are all landing (takeoff) aircraft. It follows that $\omega_{min}(\Phi_{a},Tcf_{5})=Y_{25}+Y_{53}-Y_{23}$.
(2) Suppose that aircraft $Tcf_{2}$ and $Tcf_{5}$ are both landing (takeoff) aircraft and $Tcf_{3}$ is a takeoff (landing) aircraft. It follows that $\omega_{min}(\Phi_{a},Tcf_{5})\geq\min\{T_{D}+D_{T}-Y_{13},T_{D}+D_{T}-Y_{24},0\}+Y_{25}$.
(3) Suppose that aircraft $Tcf_{3}$ and $Tcf_{5}$ are both landing (takeoff) aircraft and $Tcf_{2}$ is a takeoff (landing) aircraft. It follows that $\omega_{min}(\Phi_{a},Tcf_{5})\geq\min\{T_{D}+D_{T}-Y_{13},T_{D}+D_{T}-Y_{24},0\}+Y_{53}$.
(4) Suppose that aircraft $Tcf_{2}$ and $Tcf_{3}$ are both landing (takeoff) aircraft and $Tcf_{5}$ is a takeoff (landing) aircraft. It follows that $\omega_{min}(\Phi_{a},Tcf_{5})\geq\max\{D_{T}+T_{D}-Y_{23},0\}$.
Proof: In this proof, we only discuss the statement (2) and statements (1)(3)(4) can be proved in a similar way. When $Tcf_{4}$ is a takeoff (landing) aircraft and $Tcf_{3}$ is relevant to $Tcf_{2}$, $\omega_{min}(\Phi_{a},Tcf_{5})=Y_{25}$. When $Tcf_{4}$ is a takeoff (landing) aircraft and $Tcf_{3}$ is relevant to $Tcf_{1}$, $T_{D}+D_{T}-Y_{13}\leq 0$ and $\omega_{min}(\Phi_{a},Tcf_{5})\geq Y_{25}+T_{D}+D_{T}-Y_{13}$. That is, $\omega_{min}(\Phi_{a},Tcf_{5})\geq Y_{25}+\min\{T_{D}+D_{T}-Y_{13},0\}$. Similarly, when $Tcf_{4}$ is a landing (takeoff) aircraft, it can be proved that $S_{3}(\Phi_{b})-S_{3}(\Phi_{a})\geq\min\{T_{D}+D_{T}-Y_{13},0\}+Y_{25}$ when $Tcf_{4}$ is relevant to $Tcf_{3}$ both in $\Phi_{a}$ and $\Phi_{b}$, and $S_{4}(\Phi_{b})-S_{4}(\Phi_{a})\geq\min\{T_{D}+D_{T}-Y_{13},T_{D}+D_{T}-Y_{24},0\}+Y_{25}$ when $Tcf_{4}$ is relevant to $Tcf_{2}$ in $\Phi_{a}$ and relevant to $Tcf_{5}$ or $Tcf_{3}$ in $\Phi_{b}$. That is, $\omega_{min}(\Phi_{a},Tcf_{5})\geq\min\{T_{D}+D_{T}-Y_{13},T_{D}+D_{T}-Y_{24},0\}+Y_{25}$.
Remark 34
This theorem gives lower bounds for the time increments of the whole sequence when an aircraft is inserted, which can be used in the algorithm proposed later. Moreover, when defining $\omega_{min}$, since the aircraft $Tcf_{3}$ and $Tcf_{4}$ both have the possibilities of having no path from the last aircraft to themselves, we adopt the smaller value between $S_{3}(\Phi_{b})-S_{3}(\Phi_{a})$ and $S_{4}(\Phi_{b})-S_{4}(\Phi_{a})$ as the value of $\omega_{min}$ so as to implement a full space search for Algorithm 1 proposed later.